# Revisiting Optimal Proportions for Binary Responses: Insights from Incorporating the Absent Perspective of Type-I Error Rate Control

### v2

@@ 4.1 Neyman-like Allocation
In Section 3.4 we demonstrated that we can not simply combine use $\rho_{N_{1}}$ as an allocation rule to maximize power and test with the score test $Z_{0}$ to control the type-I error rate. To address this limitation, we propose deriving a Neyman-like allocation tailored for the $Z_{0}$ test, minimizing its variance. First, we note that the estimator $\hat{p}$ is simply the total number of successes divided by the trial size $n$, but can be expressed using the weighted average of the MLEs of each treatment arm $\hat{p}:=(n_{0}\hat{p}_{0}+n_{1}\hat{p}_{1})/n$. We can then rewrite the squared variance of the test statistic as follows
@@
$\rho^{n}_{N_{0}}=\frac{\sqrt{\hat{p}_{0}\hat{q}_{0}}}{\sqrt{\hat{p}_{0}\hat{q}_{0}}+\sqrt{\hat{p}_{1}\hat{q}_{1}}}.$ (10)
@@ 4.2 RSHIR-like Allocation
In this section we derive a proportion that minimizes the number of expected failures given a power contraint when testing with $Z_{0}$.
@@
$\text{min}_{\rho^{n}}\,\frac{\left((1-\rho^{n})(1-\hat{p}_{0})+\rho^{n}(1-\hat{p}_{1})\right)\left(\frac{\hat{p}_{0}}{\rho^{n}}+\frac{\hat{p}_{1}}{(1-\rho^{n})}-\frac{(1-\rho^{n})\hat{p}_{0}^{2}}{\rho^{n}}-2\hat{p}_{0}\hat{p}_{1}-\frac{\rho^{n}\hat{p}_{1}^{2}}{(1-\rho^{n})}\right)}{C}.$ (11)
@@
Even though we cannot find a close form solution for $\rho^{n}$ we can solve for it computationally and obtain $\rho^{n}_{R_{0}}$ (Pin, 2025). In Figure 2 we demonstrate how the proportions differ theoretically and present simulation results in the next Section.
@@
Figure 2 illustrates how the proportions change as $p_{1}$ varies, with $p_{0}$ fixed at 0.2. As indicated by the formulas, we observe that $\rho_{N_{0}}$ mirrors $\rho_{N_{1}}$. Notably, the curves for $\rho_{R_{0}}$ and $\rho_{R_{1}}$ suggest that patient allocation strategies aimed at optimizing patient outcomes differ significantly depending on the definition of the optimization problem. These differences imply that the regions where patient benefit and statistical power can be improved vary with the choice of the test statistic as well. Specifically, when the curves for $\rho_{R_{0}}$ and $\rho_{N_{0}}$ are sufficiently close, there is potential to both increase power and allocate more patients to the superior treatment arm. It is important to note that these regions shift based on the value of $p_{0}$.
@@ 4.3 Simulation Results
Building on the theoretical derivations and results, we now turn to the analysis of simulated data. As described in Section 3, we implement a moderate burn-in period of 6 patients per arm and sample with equal probability if one of the two estimated variances is zero.
Figure 1 illustrates that the two RAR designs, utilizing $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$, effectively control the type-I error rate. Notably, these methods outperform CR in terms of type-I error rate control.
In Figure 3, we present a comparison of power and the expected number of successes across the various methods. When $p_{0}=0.2$, the power of $\rho^{n}_{N_{0}}$ is comparable to that of CR. In certain regions, $\rho^{n}_{N_{0}}$ performs slightly better than CR, while in other regions, CR holds a marginal advantage. Similarly, for $p_{0}=0.5$, the power of $\rho^{n}_{N_{0}}$ and CR remains almost identical, with only minor fluctuations in performance. For a higher value of $p_{0}=0.7$, the power of both $\rho^{n}_{N_{0}}$ and CR continues to be comparable, although CR shows a slight advantage in some areas.
In the case of $\rho^{n}_{R_{0}}$, the performance is similarly close to CR at $p_{0}=0.2$, but $\rho^{n}_{N_{0}}$ outperforms CR in specific regions, and vice versa in others. However, for $p_{0}=0.5$, the power of $\rho^{n}_{R_{0}}$ is marginally inferior to CR, though the overall performance remains comparable. At $p_{0}=0.7$, $\rho^{n}_{R_{0}}$ demonstrates a more pronounced drop in power compared to CR, with the disparity being more noticeable than at $p_{0}=0.5$.
While the power of $\rho^{n}_{N_{0}}$ may be slightly better in some regions and worse in others, its overall performance in terms of ENS closely mirrors that of CR. On the other hand, $\rho^{n}_{R_{0}}$ consistently yields a higher ENS, surpassing CR by as much as 10 patients in some regions, and it never falls below the performance of CR.
Figure 3: Power and expected number of successes for complete randomization, Neyman-like $\rho^{n}_{N_{0}}$ and RSHIR-like $\rho^{n}_{R_{0}}$ testing with the score test $Z_{0}$.
Both Neyman-like $\rho^{n}_{N_{0}}$ and RSHIR-like $\rho^{n}_{R_{0}}$ proportions exhibit better control of the type-I error rate compared to CR. Neyman-like proportion achieves power levels similar to CR and in certain regions leads to an increase in the expected number of successes. In contrast, while RSHIR-like proportion may experience small power losses in some regions, it can result in significant gains in ENS, outperforming CR in this regard.
@@ 5.1 Early Phase Trial
We conducted a simulation study to evaluate the statistical properties of an early-phase, two-armed trial designed to assess the efficacy of a treatment compared to a control. Our study was based on survival outcomes observed in a multicenter randomized phase 2 trial investigating the impact of neoadjuvant chemotherapy (NAC) in patients with locally advanced esophageal cancer (EC). The trial compared patients who received two courses of NAC versus those who received three courses, revealing that a tumor reduction rate lower than 10% during the third course was associated with unfavorable overall survival (OS). Specifically, we use the 2-year OS rate of 89.3% in the more favorable group and 63.5% in the less favorable group as the basis for our assumed treatment and control arm success rates, respectively Kubo et al. (2023). The trial was designed with a total sample size of 68 patients.
@@
Procedure Type-I error Power % sup Arm (Var) ENS
$Z_{1}$ $Z_{0}$ $Z_{1}$ $Z_{0}$
@@ 5.2 Confirmatory Trial: CALISTO
In this section, we explore a real-world scenario to demonstrate the effect of the newly proposed proportions. We examine the CALISTO trial (Decousus et al., 2010). The goal of this trial was to assess the efficacy of a new drug, Arixtra, compared to a placebo in patients with acute symptomatic lower limb thrombophlebitis. Success was defined as the absence of a composite set of events, including death from any cause, symptomatic pulmonary embolism, symptomatic deep-vein thrombosis, symptomatic extension to the saphenofemoral junction, and symptomatic recurrence of superficial-vein thrombosis by day 47.
@@
Table 5 demonstrates the proportions $\rho_{CR}$, $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ effectively control the type-I error rate. Whilst, all designs have similar power both proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ are able to increase patient benefit significantly. Especially, using $\rho^{n}_{R_{0}}$ leads to 25 additional success compared to CR.
@@ 6 Discussion and Future Research
Despite the promising results, some limitations remain. We focused on two-armed trials with binary endpoints, as these are relevant in practice (Pin et al., 2025) and among the most frequently discussed settings in the theoretical literature (Hu and Rosenberger, 2006). Azriel et al. (2012) discusses the relationship to other optimal allocations for discrete and continuous endpoints that optimize for power. However, the scope of our results do not yet extend to multi-armed trials, which remain an important area for future research. The multi-armed case poses additional challenges (Tymofyeyev et al., 2007; Sverdlov et al., 2020), particularly in defining optimal allocation proportions, and requires further exploration to close this gap.
Binary endpoints, such as success or failure, are a special case in that their variance depends directly on the mean, i.e., the success rate. This influences the behavior of optimal allocation rules and differentiates binary trials from most trials with continuous outcomes. While our work addresses this specific case, optimal proportions for other endpoints will look different and do not have the same properties.
There are several promising directions for future work. First, we want to adapt our approach to alternative measures of interest, such as relative risk or odds ratios. These measures are widely used in practice and present different statistical properties that may influence optimal allocation strategies. Second, we aim to explore optimal allocation proportions for nonparametric tests and exact tests, which are particularly relevant in small-sample trials where traditional asymptotic methods may not perform well. Additionally, one could investigate deriving optimal proportions for finite sample corrections, such as the Agresti & Caffo correction, i.e. using the adjusted estimators in the optimization problem.
@@
Code
We have made our code publicly available, allowing readers to fully replicate all the results presented in this paper (Pin, 2025).
@@ Acknowledgments
The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP) and the Cusanuswerk e.V. (LP). Additionally, LP’s visit to the Department of Statistics at George Mason University to collaborate with BR on this project was supported by the Trials Methodology Research Partnership Doctoral Training Program and the Fisher Memorial Trust. LP is thankful for the departments hospitality.

### v3

@@ 4.1 Neyman-like Allocation
In Section 3.4 we demonstrated that we cannot simply use $\rho_{N_{1}}$ as an allocation rule to maximize power and test with the score test $Z_{0}$ to control the type-I error rate. To address this limitation, we propose deriving a Neyman-like allocation tailored for the $Z_{0}$ test, minimizing its variance. First, we note that the estimator $\hat{p}$ is simply the total number of successes divided by the trial size $n$, but can be expressed using the weighted average of the MLEs of each treatment arm $\hat{p}:=(n_{0}\hat{p}_{0}+n_{1}\hat{p}_{1})/n$. We can then rewrite the squared variance of the test statistic as follows
@@
$\rho^{n}_{N_{0}}=\frac{\sqrt{\hat{p}_{0}\hat{q}_{0}}}{\sqrt{\hat{p}_{0}\hat{q}_{0}}+\sqrt{\hat{p}_{1}\hat{q}_{1}}}.$
@@ 4.2 RSHIR-like Allocation
In this section we derive a proportion that minimizes the number of expected failures given a fixed variance constraint when testing with $Z_{0}$.
@@
$\text{min}_{\rho^{n}}\,\frac{\left((1-\rho^{n})(1-\hat{p}_{0})+\rho^{n}(1-\hat{p}_{1})\right)\left(\frac{\hat{p}_{0}}{\rho^{n}}+\frac{\hat{p}_{1}}{(1-\rho^{n})}-\frac{(1-\rho^{n})\hat{p}_{0}^{2}}{\rho^{n}}-2\hat{p}_{0}\hat{p}_{1}-\frac{\rho^{n}\hat{p}_{1}^{2}}{(1-\rho^{n})}\right)}{C}.$
@@
Even though we cannot find a close form solution for $\rho^{n}$ we can solve for it computationally and obtain $\rho^{n}_{R_{0}}$ (see Supplementary Material). In Figure 2 we demonstrate how the proportions differ theoretically and present simulation results in the next Section.
@@
Figure 2 illustrates how the proportions change as $p_{1}$ varies, with $p_{0}$ fixed at 0.2. As indicated by the formulas, we observe that $\rho_{N_{0}}$ mirrors $\rho_{N_{1}}$. Notably, the curves for $\rho_{R_{0}}$ and $\rho_{R_{1}}$ suggest that patient allocation strategies aimed at optimizing patient outcomes differ significantly depending on the definition of the optimization problem. These differences imply that the regions where patient-benefit and statistical power can be improved vary with the choice of the test statistic as well. Specifically, when the curves for $\rho_{R_{0}}$ and $\rho_{N_{0}}$ are sufficiently close, there is potential to both increase power and allocate more patients to the superior treatment arm. It is important to note that these regions shift based on the value of $p_{0}$.
@@ 4.3 Simulation Results
Building on the theoretical results, we now analyze the simulation for the new allocation proportions. Figure 1 shows that the score test generally yields a lower type-I error rate than the Wald test, as seen in the comparison of CR for both tests ($\rho_{CR_{0}}$ vs. $\rho_{CR_{1}}$). The two optimal optimal proportions $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$, successfully control the type-I error rate and even outperform CR in this regard.
Table 4 shows that $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$ achieve similar power to CR, with differences within $2\%$ in either direction across parameter combinations. However, the difference in type-I error rate control is substantial: CR fails to control the nominal level, while both new proportions do. This does demonstrate that neither of the two designs using the new allocation proportion requires any of the potential solutions discussed in Section 3, including larger burn-in sizes. While $\rho^{n}_{N_{0}}$ may reduce the ENS compared to CR, $\rho^{n}_{R_{0}}$ tends to increase it and typically allocates more patients to the superior arm.
@@ 5.1 Early Phase Trial
We redesign a study to evaluate the statistical properties of an early-phase, two-armed trial designed to assess the efficacy of a treatment compared to a control. Our study was based on survival outcomes observed in a multicenter randomized phase 2 trial investigating the impact of neoadjuvant chemotherapy (NAC) in patients with locally advanced esophageal cancer (EC). The trial compared patients who received two courses of NAC versus those who received three courses, revealing that a tumor reduction rate lower than 10% during the third course was associated with unfavorable overall survival (OS). Specifically, we use the 2-year OS rate of 89.3% in the more favorable group and 63.5% in the less favorable group as the basis for our assumed treatment and control arm success rates, respectively Kubo et al. (2023). The trial was designed with a total sample size of 68 patients.
@@
Test Procedure Type-I error Power % sup Arm (Var) ENS
@@ 5.2 Confirmatory Trial: CALISTO
In this section, we redesign a published clinical trial to demonstrate the effect of the newly proposed proportions. We examine the CALISTO trial (Decousus et al., 2010). The goal of this trial was to assess the efficacy of a new drug, Arixtra, compared to a placebo in patients with acute symptomatic lower limb thrombophlebitis. Success was defined as the absence of a composite set of events, including death from any cause, symptomatic pulmonary embolism, symptomatic deep-vein thrombosis, symptomatic extension to the saphenofemoral junction, and symptomatic recurrence of superficial-vein thrombosis by day 47.
@@
Table 5 demonstrates the proportions $\rho_{CR}$, $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ effectively control the type-I error rate. While all designs have similar power, both proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ are able to increase patient benefit significantly. Especially, using $\rho^{n}_{R_{0}}$ leads to 25 additional success compared to CR.
@@ 6 Discussion and Future Research
Despite the promising results, some limitations remain. We focused on two-armed trials with binary endpoints, as these are relevant in practice (Pin et al., 2025) and are among the most frequently discussed settings in the theoretical literature (Hu and Rosenberger, 2006). Azriel et al. (2012) discusses the relationship to other optimal allocations for discrete and continuous endpoints that optimize for power. However, the scope of our results do not yet extend to multi-armed trials, which remain an important area for future research.
While our focus in this paper is on two-arm trials, there is a large literature on optimal allocation proportions in multi-armed settings. However, the multi-armed case poses additional challenges (Tymofyeyev et al., 2007; Sverdlov et al., 2020). For instance, multi-armed extensions of Neyman allocation exist (Tymofyeyev et al., 2007), though they tend to allocate patients only to the best and worst arms, which is undesirable in practice. As shown in Zhu and Hu (2009); Baldi Antognini et al. (2021), additional constraints (e.g., minimum allocation proportions) are needed to ensure all arms are represented. Other authors have proposed allocations tailored to specific hypothesis tests or altered objectives (Jeon and Hu, 2010; Azriel and Feigin, 2014; Biswas et al., 2011). Our proposed proportions, which are optimized specifically for the score test in the two-arm setting, likely extend in the same way the multi-arm case but also have the same limitations. While a score-test analog of the multi-armed Neyman allocation could, in principle, be derived, it would also only allocate to two arms unless further constraints are imposed. Tymofyeyev et al. (2007) investigated multi-armed extensions of both Neyman and RSHIR allocation, but the latter solution requires the choice of a smoothing kernel function and the choice of unknown parameters, but do not provide a closed-form solution.
An alternative route of approaching type-I error rate control for optimal proportions would be to incorporate it directly as a constraint in the optimization problem. This approach has been discussed, for example in Baas et al. (2024), but this would require separate computation for each sample size $n$ over the entire parameter space. Instead, we aim to highlight the practical benefits of using better suited test statistics within standard adaptive design frameworks.
Lastly, binary endpoints, such as success or failure, are a special case in that their variance depends directly on the mean, i.e., the success rate. This influences the behavior of optimal allocation rules and differentiates binary trials from most trials with continuous outcomes. While our work addresses this specific case, optimal proportions for other endpoints will look different and do not have the same properties. Ye et al. (2024) explored the robustness of RAR and the use of the Wald test for continuous responses and large samples.
There are several promising directions for future work. First, we want to adapt our approach to alternative measures of interest, such as relative risk or odds ratios. The calculations are straightforward for the Wald test (Rosenberger et al., 2001; Pin et al., 2024) and can be equivalently derived for the score test. However, these measures are widely used in practice and present different statistical properties that may influence optimal allocation strategies. Second, we aim to explore optimal allocation proportions for nonparametric tests and exact tests, which are particularly relevant in small-sample trials where traditional asymptotic methods may not perform well. Additionally, one could investigate deriving optimal proportions for finite sample corrections, such as the Agresti and Caffo correction, i.e. using the adjusted estimators in the optimization problem.
@@ Acknowledgments
The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP) and the Cusanuswerk e.V. (LP). SSV is part of PhaseV’s advisory board. Additionally, LP’s visit to the Department of Statistics at George Mason University to collaborate with WR on this project was supported by the Trials Methodology Research Partnership Doctoral Training Program and the Fisher Memorial Trust. LP is thankful for the departments hospitality.
Supplementary Material (Code)
The code referenced in Section 4.2 is available with this paper at the Biometrics website on Oxford Academic. The code can be found at the GitHub repository https://github.com/lukaspinpin/Optimal-Proportions, and contains the R code necessary to replicate the results presented in the paper.
Data Availability
There are no data presented in this paper.